OpenStudy (anonymous):

how do i subtract 1/9 - 1/6?

OpenStudy (anonymous):

Firs take the LCD of 9and 6

OpenStudy (anonymous):

Find the least common denominator (18) and subtract 2/18 from 3/18.

OpenStudy (electrochika):

2/18 - 3/18

OpenStudy (anonymous):

LCD of 9,6 is 18 1/9=2/18 1/6=3/18

OpenStudy (anonymous):

ok thanks!

OpenStudy (anonymous):

1/9-1/6=2/18-3/18 (you can subtract only when you have common denominator) 2-3=-1 -1/18 should be your answer